(x-y)p-(x-y+z)q=z


1
Expert's answer
2021-05-02T16:01:24-0400

The equation

"(x-y)p-(x-y+z)q=z"

is a Lagrange linear PDE of the form:


"Pp+Qq=R"

where:

"P=x-y\\ Q=-(x-y+z)\\
R= z"


The auxiliary equation is of the form


"\frac{dx}{P}=\frac{dy}{Q}=\frac{dz}{z}"

Thus, from the given equation, we have:


"\frac{dx}{x-y}= \frac{dy}{-(x-y+z)}=\frac{dz}{z} \qquad \cdots (i)"

Solving equation (i) we get:


"\frac{dx+dy+dz}{x-y+[-(x-y+z)]+z} = k \qquad \text{(say)}\\
\frac{dx+dy+dz}{x-y-(x-y+z)+z} = k\\
\frac{dx+dy+dz}{x-y-x+y-z+z} = k\\
\frac{dx+dy+dz}{0} = k\\
\implies dx+dy+dz =0"

Integrating through, we get:


"\int dx+\int dy+\int dz = \int 0\\
x+y+z= c"

this implies that:


"\bold{c_1 = x+y+z}"

From equation (i), we can have:


"\frac{dx+dy}{x-y-(x-y-z)}=\frac{dz}{z}\\
\frac{dx+dy}{x-y-x+y+z}=\frac{dz}{z}\\
\frac{dx+dy}{z}=\frac{dz}{z}\\
\implies dx+dy=dz"

Integrating through, we get:


"\int dx + \int dy = \int dz\\
x+y=z+c_2"

This implies that:


"\bold{c_2 = x+y-z}"


The general solution is thus:


"\phi(c_1,c_2) =0 \\\implies \phi(x+y+z\;,\quad x+y-z) =0"